# Gaussian rational numbers in Cantor sets in the complex plane

### v1

@@ 1. Introduction
Closely related to Mahler’s problem is the question of what the rational numbers in the middle-third Cantor set are. Wall [29] first noted that $\frac{1}{4}$ and $\frac{3}{4}$ are the only dyadic rationals in $\bm{C}$. He also showed in [30] that there are exactly $14$ rationals in $\bm{C}$ with decimal denominators. Later, Nagy [20] proved that for any prime number $p>3$, there are only finitely many rational numbers in $\bm{C}$ whose denominators are powers of $p$. Bloshchitsyn [3] obtained analogous results for the class of $m$-adic Cantor sets, a generalization of $\bm{C}$ obtained by replacing the base $3$ with an integer $m\geq 3$ and the digit set $\{0,2\}$ with a proper subset of $\{0,1,\ldots,m-1\}$ in (1.1). Recently, Schleischitz [24] proved a more general result in this setting: if $P$ is a finite set of distinct primes not dividing $m$, then there are only finitely many rational numbers in such a set whose denominators have all prime factors from $P$. Shparlinski [26] later obtained a quantitative version of this result using a different method. Li et al. [16] extended Shparlinski’s result to the broader setting of $\times m$-invariant sets on $[0,1]$. Jiang et al. [11] proved a finiteness result for translations of $m$-adic Cantor sets. Very recently, Kong et al. [13] obtained a further extension of Schleischitz’s result by allowing the digit set to be a finite subset of $\mathbb{Q}$.
Solomon, Trauthwein and Weiss [22] studied the problem of counting rational numbers in $\bm{C}$ up to a given height. They formulated a conjecture asserting that there are $\ll N^{\log_{3}2}$ rational numbers in $\bm{C}$ with denominators at most $N$. This conjecture (and its natural generalizations) remains open; see [5, 6] for significant progress. Here and throughout, the notation $a(x)\ll b(x)$ means that $a(x)\leq cb(x)$ for some positive constant $c$ independent of $x$.
@@ 3. Period estimate for Gaussian rational numbers in $K_{\beta,D}$
In this section, we apply the counting result (Theorem 2.5) to give an upper bound estimate for the period length of an eventually period coding for every Gaussian rational number in $K_{\beta,D}$.
@@ 5. Number-theoretical preliminaries
Let $\alpha\in\mathbb{Z}[i]$ and let $\gamma\in\mathbb{Z}[i]$ be a prime element with $\gamma\nmid\alpha$. Set $d={\rm ord}(\alpha;\gamma)$ and $m=\nu_{\gamma}(\alpha^{d}-1)$. Then
${\rm ord}(\alpha;\gamma)={\rm ord}(\alpha;\gamma^{2})=\cdots={\rm ord}(\alpha;\gamma^{m})=d.$ (5.1)
Moreover, if $\gamma$ is of Type I or Type III, then
${\rm ord}(\alpha;\gamma^{m+j})=p^{j}d,\quad\nu_{\gamma}(\alpha^{p^{j}d}-1)=m+j,\quad\forall j\in\mathbb{N},$ (5.2)
where $p=H(\gamma)$. If $\gamma$ is of Type II, then
${\rm ord}(\alpha;\gamma^{m+2j})=2^{j},\quad\nu_{\gamma}(\alpha^{2^{j}}-1)=m+2j,\quad\forall j\in\mathbb{N}.$ (5.3)
@@
Since ${\rm ord}(\alpha;\gamma)=d$ and $\nu_{\gamma}(\alpha^{d}-1)=m$, we have for $1\leq k\leq m$, $\alpha^{d}\equiv 1\pmod{\gamma^{k}}$ and so ${\rm ord}(\alpha;\gamma^{k})\mid d$. On the other hand, by Lemma 5.1 we have $d\mid{\rm ord}(\alpha;\gamma^{k})$. Hence ${\rm ord}(\alpha;\gamma^{k})=d$ for $1\leq k\leq m$. This proves (5.1). Below we prove (5.2) by induction on $j$.
Suppose $\gamma$ is of either Type I or Type III. Since $\nu_{\gamma}(\alpha^{d}-1)=m$, we can write $\alpha^{d}=1+\omega\gamma^{m}$, where $\omega\in\mathbb{Z}[i]$, $\gamma\nmid\omega$. Then by the binomial formula,
@@
By Lemma 5.4, $\gamma\mid p$ and $\nu_{\gamma}(p)=1$. It follows that $\alpha^{pd}\equiv 1\pmod{\gamma^{m+1}}$ and $\nu_{\gamma}(\alpha^{dp}-1)=m+1$. Hence ${\rm ord}(\alpha;\gamma^{m+1})\mid pd$. On the other hand, by Lemma 5.1, $d\mid{\rm ord}(\alpha;\gamma^{m+1})$. Moreover, $\nu_{\gamma}(\alpha^{d}-1)=m$ implies that ${\rm ord}(\alpha;\gamma^{m+1})\neq d$. Hence ${\rm ord}(\alpha;\gamma^{m+1})=pd$. It then follows from (5.4) that $\nu_{\gamma}(\alpha^{pd}-1)=m+1$. This proves (5.2) for $j=1$.
Suppose that (5.2) holds for some $j\geq 1$. We are going to prove that it also holds for $j+1$. To this end, note that by the induction hypothesis, we can write $\alpha^{p^{j}d}=1+\eta\gamma^{m+j}$, where $\eta\in\mathbb{Z}[i]$ and $\gamma\nmid\eta$. Then by applying the binomial formula to $\alpha^{p^{j+1}d}=(1+\eta\gamma^{m+j})^{p}$ and a similar argument as above, we see that (5.2) holds for $j+1$. This proves (5.2).
To see (5.3), let $\gamma$ be a Type II prime element. Since $\mathbb{Z}[i]$ is a Euclidean domain with the norm NN, $\gamma\nmid\alpha$ and $N(\gamma)=2$, we easily see from the division algorithm that $\alpha\equiv 1\pmod{\gamma}$. Hence $d=1$. Then (5.3) follows by applying Lemma 5.4 and a similar argument as above. ∎
@@
We have shown in the proof of Lemma 5.7 that $\alpha\equiv 1\pmod{\gamma}$. Let $m=\nu_{\gamma}(\alpha-1)$. For $n\geq m$, let $j\geq 0$ be such that
@@
Then by Lemma 5.1 and Lemma 5.7, ${\rm ord}(\alpha;\gamma^{n})$ is an integer power of $2$ and
@@
$2,\,H(\eta_{1}),\,\ldots\,H(\eta_{q}),\,H(\beta_{1}),\,\ldots,\,H(\beta_{k})$
@@
Then by (5)-(5.5), the assumption that $H(\eta_{i})=H(\xi_{i})$ for $1\leq i\leq q$, Lemmas  5.1 and 5.8, we see that
@@
Now we are ready to give the proof of Theorem 1.1. The core of the proof is the application of Lemma 3.1 and Lemma 5.9, together with an argument from [13].
@@
Acknowledgements. The author would like to thank Dejun Feng, Jiangtao Li, Ruofan Li, Caiyun Ma and Zhao Shen for helpful comments and suggestions. This project is supported by the Natural Science Foundation of China (Grant No. 12301110) and Natural Science Foundation of Hunan Province (Grant No. 2023JJ40700).

### v2

@@ 1. Introduction
Closely related to Mahler’s problem is the question of what the rational numbers in the middle-third Cantor set are. Wall [29] first noted that $0,\frac{1}{4},\frac{3}{4}$ and $1$ are the only dyadic rationals in $\bm{C}$. He also showed in [30] that there are exactly $14$ rationals in $\bm{C}$ with decimal denominators. Later, Nagy [20] proved that for any prime number $p>3$, there are only finitely many rational numbers in $\bm{C}$ whose denominators are powers of $p$. Bloshchitsyn [3] obtained analogous results for the class of $m$-adic Cantor sets, a generalization of $\bm{C}$ obtained by replacing the base $3$ with an integer $m\geq 3$ and the digit set $\{0,2\}$ with a proper subset of $\{0,1,\ldots,m-1\}$ in (1.1). Recently, Schleischitz [24] proved a more general result in this setting: if $P$ is a finite set of distinct primes not dividing $m$, then there are only finitely many rational numbers in such a set whose denominators have all prime factors from $P$. Shparlinski [26] later obtained a quantitative version of this result using a different method. Li et al. [16] extended Shparlinski’s result to the broader setting of $\times m$-invariant sets on $[0,1]$. Jiang et al. [11] proved a finiteness result for translations of $m$-adic Cantor sets. Very recently, Kong et al. [13] obtained a further extension of Schleischitz’s result by allowing the digit set to be a finite subset of $\mathbb{Q}$.
Solomon, Trauthwein and Weiss [22] studied the problem of counting rational numbers in $\bm{C}$ up to a given height. They formulated a conjecture asserting that for any $\epsilon>0$, there are $\ll N^{\log_{3}2+\epsilon}$ reduced rational numbers in $\bm{C}$ with denominators at most $N$. This conjecture (and its natural generalizations) remains open; see [5, 6] for significant progress. Here and throughout, the notation $a(x)\ll b(x)$ means that $a(x)\leq cb(x)$ for some positive constant $c$ independent of $x$.
@@ 3. Period estimate for Gaussian rational numbers in $K_{\beta,D}$
In this section, we apply the counting result (Theorem 2.5) to give an upper bound estimate for the period length of an eventually periodic coding for every Gaussian rational number in $K_{\beta,D}$.
@@ 5. Number-theoretical preliminaries
Let $\alpha,\gamma\in\mathbb{Z}[i]$ with $\gamma$ a prime element and $\gamma\nmid\alpha$. If $\gamma$ is of Type I or Type III, set $p=H(\gamma)$, $d={\rm ord}(\alpha;\gamma)$ and $m=\nu_{\gamma}(\alpha^{d}-1)$. Then we have
${\rm ord}(\alpha;\gamma)={\rm ord}(\alpha;\gamma^{2})=\cdots={\rm ord}(\alpha;\gamma^{m})=d,$ (5.1)
and
${\rm ord}(\alpha;\gamma^{m+j})=p^{j}d,\quad\nu_{\gamma}(\alpha^{p^{j}d}-1)=m+j,\quad\forall j\in\mathbb{N}.$ (5.2)
If $\gamma$ is of Type II, let $u\in\mathbb{Z}[i]$ be the unit such that $2=u\gamma^{2}$. Set $m=\nu_{\gamma}(\alpha-1)$ and $t_{j}=\nu_{\gamma}(\alpha^{2^{j}}-1)$ for $j\in\mathbb{N}$. Then the following hold.
(i) If $m=1$, set $\tau_{1}=\nu_{\gamma}(\eta^{2}+u\gamma\eta+u)$, where $\eta=(\alpha-1)\gamma^{-1}$. Then
$t_{1}=2,\quad t_{j}=\tau_{1}+2j\quad(\forall j\geq 2),\quad{\rm ord}(\alpha;\gamma^{t_{j}})=2^{j}\quad(\forall j\in\mathbb{N}).$ (5.3)
(ii) If $m=2$, set $\tau_{2}=\nu_{\gamma}(\eta+u)$, where $\eta=(\alpha-1)\gamma^{-2}$. Then
$t_{j}=\tau_{2}+2j+2,\quad{\rm ord}(\alpha;\gamma^{t_{j}})=2^{j},\quad\forall j\in\mathbb{N}.$
(iii) If $m\geq 3$, then $t_{j}=m+2j$ and ${\rm ord}(\alpha;\gamma^{t_{j}})=2^{j}$ for $j\in\mathbb{N}$.
@@
First consider the case that $\gamma$ is of either Type I or Type III. Since ${\rm ord}(\alpha;\gamma)=d$ and $\nu_{\gamma}(\alpha^{d}-1)=m$, we have for $1\leq k\leq m$, $\alpha^{d}\equiv 1\pmod{\gamma^{k}}$ and so ${\rm ord}(\alpha;\gamma^{k})\mid d$. On the other hand, by Lemma 5.1 we have $d\mid{\rm ord}(\alpha;\gamma^{k})$. Hence ${\rm ord}(\alpha;\gamma^{k})=d$ for $1\leq k\leq m$. This proves (5.1). Below we prove (5.2) by induction on $j$.
Since $\nu_{\gamma}(\alpha^{d}-1)=m$, we can write $\alpha^{d}=1+\omega\gamma^{m}$, where $\omega\in\mathbb{Z}[i]$, $\gamma\nmid\omega$. Then by the binomial formula,
@@
By Lemma 5.4, $\gamma\mid p$ and $\nu_{\gamma}(p)=1$. It follows that $\alpha^{pd}\equiv 1\pmod{\gamma^{m+1}}$ and $\nu_{\gamma}(\alpha^{dp}-1)=m+1$. Hence ${\rm ord}(\alpha;\gamma^{m+1})\mid pd$. On the other hand, by Lemma 5.1, $d\mid{\rm ord}(\alpha;\gamma^{m+1})$. Moreover, $\nu_{\gamma}(\alpha^{d}-1)=m$ implies that ${\rm ord}(\alpha;\gamma^{m+1})\neq d$. Hence ${\rm ord}(\alpha;\gamma^{m+1})=pd$. It then follows from (5.4) that $\nu_{\gamma}(\alpha^{pd}-1)=m+1$. This proves (5.2) for $j=1$. Suppose that (5.2) holds for some $j\geq 1$. We are going to prove that it also holds for $j+1$. To this end, note that by the induction hypothesis, we can write $\alpha^{p^{j}d}=1+\xi\gamma^{m+j}$, where $\xi\in\mathbb{Z}[i]$ and $\gamma\nmid\xi$. Then again by the binomial formula, we have
$\alpha^{p^{j+1}d}=(1+\xi\gamma^{m+j})^{p}=1+p\xi\gamma^{m+j}+\binom{p}{2}(\xi\gamma^{m+j})^{2}+\cdots+\binom{p}{p}(\xi\gamma^{m+j})^{p}.$
Again, since $\gamma\mid p$ and $\nu_{\gamma}(p)=1$, we see that $\nu_{\gamma}(\alpha^{p^{j+1}d}-1)=m+j+1$. Hence ${\rm ord}(\alpha;\gamma^{m+j+1})\mid p^{j+1}d$. It then follows from Lemma 5.1 and the induction hypothesis that ${\rm ord}(\alpha;\gamma^{m+j+1})=p^{j+1}d$. Therefore, (5.2) holds for $j+1$. This proves (5.2).
Next, we consider the case that $\gamma$ is of Type II. Since $\mathbb{Z}[i]$ is a Euclidean domain with the norm $N$, $\gamma\nmid\alpha$ and $N(\gamma)=2$, we easily see from the division algorithm that $\alpha\equiv 1\pmod{\gamma}$. Hence $m=\nu_{\gamma}(\alpha-1)$ is well-defined. Below we only present the proof for the case that $m=1$; the other two cases can be verified similarly.
Since $m=1$, we have $\alpha=1+\gamma\eta$ with $\gamma\nmid\eta$. Then
$\alpha^{2}=(1+\gamma\eta)^{2}=1+2\gamma\eta+\gamma^{2}\eta^{2}=1+\gamma^{2}\eta(u\gamma+\eta).$
Since $\gamma\nmid\eta$ and $\gamma\nmid(u\gamma+\eta)$, it follows that $\nu_{\gamma}(\alpha^{2}-1)=2$ and ${\rm ord}(\alpha;\gamma^{2})\mid 2$. Since $m=1$, we have ${\rm ord}(\alpha;\gamma^{2})=2$. For $\alpha^{2^{2}}$ we have
$\alpha^{2^{2}}$ $=(1+\gamma^{2}\eta(u\gamma+\eta))^{2}$
$=1+2\gamma^{2}\eta(u\gamma+\eta)+\gamma^{4}\eta^{2}(u\gamma+\eta)^{2}.$
$=1+\gamma^{4}\eta(u\gamma+\eta)(\eta^{2}+u\gamma\eta+u).$
Since $\tau_{1}=\nu_{\gamma}(\eta^{2}+u\gamma\eta+u)$, it follows that $\nu_{\gamma}(\alpha^{2^{2}}-1)=\tau_{1}+4$. This proves that the statement for $t_{j}$ in (5.3) holds for $j=2$. Then by induction on $j$, using the relation $2=u\gamma^{2}$ and arguing as in the Type I/III case, we obtain the formulas for $t_{j}$ and ${\rm ord}(\alpha;\gamma^{t_{j}})$ for all $j\geq 1$ in (5.3). This finishes the proof of the lemma. ∎
@@
We have seen in the proof of Lemma 5.7 that $\alpha\equiv 1\pmod{\gamma}$. Let $m=\nu_{\gamma}(\alpha-1)$. For simplicity, we only consider the case that $m\geq 3$; the other two cases $m=1$ and $m=2$ can be proved similarly with the aid of Lemma 5.7 (with possibly different constants $C_{2}$). For $n\geq m$, let $j\geq 0$ be such that
@@
Then by Lemma 5.1 and Lemma 5.7(iii), ${\rm ord}(\alpha;\gamma^{n})$ is an integer power of $2$ and
@@
$2,\,H(\eta_{1}),\,\ldots\,H(\eta_{q}),\,H(\beta_{1}),\,\ldots,\,H(\beta_{k})$ (5.5)
@@
Then by (5.5)-(5.6), the assumption that $H(\eta_{i})=H(\xi_{i})$ for $1\leq i\leq q$, Lemmas  5.1 and 5.8, we see that
@@
Now we are ready to give the proof of Theorem 1.1. The core of the proof is the application of Lemmas 3.1-5.9, together with an argument from [13].
@@
Acknowledgements. The author would like to thank De-Jun Feng, Bing Li, Jiangtao Li, Ruofan Li, Cai-Yun Ma and Zhao Shen for their helpful comments and suggestions. This project is supported by the Natural Science Foundation of China (Grant No. 12301110) and Natural Science Foundation of Hunan Province (Grant No. 2023JJ40700).